OpenStudy (anonymous):

Simplify as much as possible. 8w sqrt(3x^3) - x sqrt(75xw^2)

OpenStudy (anonymous):

Help please!!

OpenStudy (anonymous):

\[8wx \sqrt{3x} - 5wx \sqrt{3x}\]

OpenStudy (anonymous):

@kaka2266 can u find the answer now?

OpenStudy (anonymous):

so should i simplify it more?

OpenStudy (anonymous):

yes just subtract 2 terms

OpenStudy (anonymous):

Okey one moment i will try it

OpenStudy (anonymous):

3wx\[\sqrt{3x}\]

OpenStudy (anonymous):

is that correct ?

OpenStudy (anonymous):

exactly u r right

OpenStudy (anonymous):

\[=3wx \sqrt{3x}\]

OpenStudy (anonymous):

Thank u so much

OpenStudy (anonymous):

ur welcome